The Marine Division of Pacific Corp. has average invested assets of $100,000,000. Sales revenue of $50,260,000 results in an operating income of $9,973,000. The hurdle rate is 8%.

a. Calculate the return on investment. (Round your answer to 2 decimal places.)



b. Calculate the profit margin. (Round your answer to 2 decimal places.)



c. Calculate the investment turnover. (Round your answer to 4 decimal places.)



d. Calculate the residual income.
Answer #1

a.ROI=Net operating income/Average operating assets

=(9,973,000/100,000,000)=9.97%(Approx).

b.Profit margin=Net operating income/Sales

=(9,973,000/50,260,000)=19.84%(Approx).

c.Investment turnover=Sales/Average operating assets

=(50,260,000/100,000,000)=0.5026

d.Residual income=Net operating income-(Average operating assets*Hurdle Rate)

=9,973,000-(100,000,000*8%)

=1,973,000